OpenStudy (moonlitfate):

The population P of a bacteria culture is modeled by P = 4100e^(k*t), where t is the time in hours. If the population of the culture was 5800 after 40 hours, how long does it take for the population to double? Round to the nearest tenth of an hour.

OpenStudy (helder_edwin):

u have \[ \large P=4100e^{kt}. \] for t=40 u have P=5800 then \[ \large 5800=4100e^{40k} \] \[ \large 1.4146=\frac{58}{41}=\frac{5800}{4100}=e^{40k} \] \[ \large \ln(1.4146)=40k \] \[ \large k=\frac{40}{\ln(1.4146)}=115.3167 \]

OpenStudy (helder_edwin):

therefore \[ \large P=4100e^{115.3167\times t} \]

OpenStudy (helder_edwin):

finally \[ \large 11600=5800\times2=4100e^{115.3167\times t} \] \[ \large 2.8293=\frac{11600}{4100}=e^{115.3167\times t} \] \[ \large \ln(2.8293)=115.3167t \] \[ \large t=\frac{115.3167}{\ln(2.8293)}=110.8795 \] so in approximately 111 hours, the population will double.

OpenStudy (moonlitfate):

110 isn't one of the available answers. And I've tried re-working the problem and nothing really seems to work. Frustrating.

OpenStudy (lgbasallote):

slight mistake in solving for k @helder_edwin

OpenStudy (helder_edwin):

my bad sorry \[ \large k=\frac{\ln(1.4146)}{40}=0.0087 \]

OpenStudy (helder_edwin):

yeah i saw it

OpenStudy (lgbasallote):

better

OpenStudy (moonlitfate):

I managed to solve for k; k = 0.0087 ; not sure what to do then. x:

OpenStudy (helder_edwin):

therefore \[ \large P=4100e^{0.0087t} \]

OpenStudy (lgbasallote):

you can use a shortcut actually \[\huge t = \frac{\ln 2}{k}\] where t is the time it takes for the bacteria to double

OpenStudy (helder_edwin):

so \[ \large t=\frac{\ln(2.8293)}{0.0087}=119.9314 \]

OpenStudy (moonlitfate):

Hm. That was the same answer I would get when I was trying to solve the problem, but that was wrong, too. :/

OpenStudy (moonlitfate):

Answers are: A) 56.6 hours B) 54.8 hours C) 81.7 hours D) 79.9 hours E)8.9 hours

OpenStudy (lgbasallote):

try using this \[t = \frac{ln 2}{0.00867}\]

OpenStudy (lgbasallote):

you'll get one of those choices

OpenStudy (helder_edwin):

maybe it is asking for the time it will take for the population to double from its starting point.

OpenStudy (lgbasallote):

you were doubling the given value?

OpenStudy (lgbasallote):

if you're going to double 5800 then why go through all the trouble of using logarithms...just double 40 hours and add 40....

OpenStudy (helder_edwin):

so \[ \large 8200=4100e^{0.0087t} \] \[ \large \ln(2)=0.0087t \] \[ \large t=\frac{\ln(2)}{0.0087}=79.6721 \]

OpenStudy (moonlitfate):

Just confused right now. :(

OpenStudy (lgbasallote):

@helder_edwin use 0.00867 as the constant...

OpenStudy (helder_edwin):

in that case t=79.94777

OpenStudy (lgbasallote):

yes. it shows the choices more

OpenStudy (moonlitfate):

Thanks guys. I'll try to remember that. .:)